A private plain traveled from Seattle to the rugged wilderness, at an average speed of 156 mph.on the return trip the average speed was 182 mph if the total traveling time was 9 hours, how far was Seattle from the wilderness?
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

let distance be x miles
time going + time return = 9
x/156 + x/182=9
182x+156x= 9 *182*156
338x=25528
x=756 miles

Answer by Alan3354(69443)

A private plain traveled from Seattle to the rugged wilderness, at an average speed of 156 mph.on the return trip the average speed was 182 mph if the total traveling time was 9 hours, how far was Seattle from the wilderness?
--------------------
Avg speed for the round trip = 2*156*182/(156+182)
= 168 mi/hr
-----------
RT distance = 168*9 = 1512 miles
One way = 756 miles
========================
Notice the formula for avg speed is similar to parallel resistors and parallel flows in pipes, with a factor of 2 in the numerator.
------------------
The distance is not a factor.